2a+b/b=3a-1/6 solve for a
Answer by stanbon(75887)

(2a+b)/b = (3a-1)/6 solve for a
--------
Cross-Multiply to get:
12a + 6b = b(3a-1)
----
12a + 6b = 3ab - b
------
a(12-3b) = -7b
------
a = -7b/(12-3b)
-------------------
Cheers,
Stan H.
===================